A 16 Ohm and a 20 Ohm resistor are connected in parallel. A difference of potential of 40 V is applied to the combination. a) Compute the equivalent resistance of the parallel circuit. b) What is the current in the circuit? c)How large is the current through the 16 Ohm resistor?


1
Expert's answer
2021-05-21T10:55:35-0400

a)

"\dfrac{1}{R_{eq}}=\dfrac{1}{R_1}+\dfrac{1}{R_2},""R_{eq}=\dfrac{1}{\dfrac{1}{R_1}+\dfrac{1}{R_2}}=\dfrac{1}{\dfrac{1}{16\ \Omega}+\dfrac{1}{20\ \Omega}}=8.9\ \Omega."

b)

"I=\dfrac{V}{R_{eq}}=\dfrac{40\ V}{8.9\ \Omega}=4.5\ A."

c)

"I_1=\dfrac{V_1}{R_1}=\dfrac{40\ V}{16\ \Omega}=2.5\ A."